# Estimated Heating Rates Due to Cyclotron Damping of Ion-scale Waves Observed by Parker Solar Probe

### v2

@@ 2.1 Plasma and Magnetic Field Data
We use Level 2 magnetic field data (in spacecraft frame RTN coordinates) from the Flux Gate Magnetometer (MAG) component of the PSP/FIELDS instrument suite, which makes in situ observations of electromagnetic fields (Bale et al., 2016). While the sampling rate of the MAG varies between $\sim$73.24 Sa $s^{-1}$, 146.48 Sa $s^{-1}$, and 292.96 Sa $s^{-1}$ during Encounter 1, the magnetic field is sampled at a constant rate of 146.48 Sa $s^{-1}$ during Encounter 2. Measurements of the thermal plasma come from the PSP/SWEAP instrument suite (Kasper et al., 2016). Huang et al. (2020) and Halekas et al. (2020) have estimated parallel and perpendicular temperatures (with respect to the direction of magnetic field) of protons ($T_{\parallel,p}$ and $T_{\perp,p}$) and electrons ($T_{\parallel,e}$ and $T_{\perp,e}$), respectively by applying bi-Maxwellian fits to velocity distribution functions (VDFs) of protons evaluated from SWEAP/SPC (Case et al., 2020) and electrons evaluated from SWEAP/SPANe (Whittlesey et al., 2020) measurements. We use plasma parameters evaluated by averaging over 1 minute of SPC and SPANe observations as inputs into the PLUME solutions.
@@
Figure 2: Overview of methodology to estimate ion-scale wave dissipation in the PSP interval (considered in Figure 1) on 10th November 2018 17:53:58 - 18:08:53 where both RH and LH polarized waves are observed: (a) The spacecraft-frame polarization spectrum, $\sigma(\omega,t)$, (b) the frequency ranges where LH and RH polarized waves are observed ($\omega_{\text{obs}}^{LH/RH}$, red and blue dash-dotted, respectively), and (c) their corresponding energy spectra (red and blue) along with the total energy spectrum (black). (d, e) The damping rates of ICWs and FMWs onto protons and electrons, respectively as functions of spacecraft-frame frequencies that are evaluated using PLUME for all possible Doppler shifts.
@@ 2.3 Removing non-wave Signatures and Identifying Ion-scale Wave Frequencies in the Spacecraft Frame
In SKM23, we observed significant energy in ion-scale waves in the first two PSP encounters, as has been previously reported, e.g. Bowen et al. (2020b), Verniero et al. (2020), Zhao et al. (2021), and Liu et al. (2023). Panel (a) of Figure 2 shows $\sigma(s,t)$ for the same example PSP 15-minute interval on November 10, 2018, where ion-scale waves are observed. Persistent high values of $|\sigma|$ associated with these waves are observed over frequency bands near the proton cyclotron frequency, $f_{cp}=\frac{q_{p}B}{2\pi m_{p}}$, and last for times varying from several seconds to a few hours. Here $q_{p}$ is the charge of the proton. In the absence of ion-scale waves, non-persistent random $\sigma$ values corresponding to the background turbulent fluctuations are observed over timescales of $\sim$ 1/frequency, i.e. the time resolution of the wavelet transform. By employing a time-smoothing routine on $\sigma(s,t)$, we identify the frequency band, $f^{(RH/LH)}_{\text{obs}}$ where the ion-scale waves are observed (shown in panel (b) of Figure 2) and evaluate their energy spectra, $|\tilde{B}(f)|^{2^{LH/RH}}_{wave\_psp}$ (shown in panel (c) of Figure 2) as described in Appendix A. Here $\omega^{(RH/LH)}_{\text{obs}}=2\pi f^{(RH/LH)}_{\text{obs}}$ is the angular frequency band where the waves are observed.
@@ 2.4 Calculating Damping Rates at $\omega_{\text{obs}}$
The four MHD modes - the Alfvén, fast and slow magnetosonic waves, and the entropy mode have kinetic counterparts in a collisionless plasma that can be relatively lightly damped. The assumption of symmetric VDFs produces degenerate solutions for the three propagating modes, and therefore the damping rates of the forward and backwards solutions are identical. The impact of the mode propagation direction is considered in the calculation of the spacecraft to plasma frame conversion, see Eqn. 2. At small parallel scales, Alfvén waves extend to ICWs and fast magnetosonic waves extend to whistler waves. The PLUME dispersion solver (Klein & Howes, 2015) calculates the hot linear plasma dispersion relation $|D(\omega,\gamma;\mathbf{k},\mathcal{P})|=0$ for bi-Maxwellian VDFs of ions and electrons in a uniform magnetized environment. The solutions of the dispersion relation yield dimensionless angular frequencies and damping rates $(\omega/\Omega_{p},\gamma/\Omega_{p})_{\texttt{PLUME}}$ in the plasma frame–which are normalized to physical units via the observed proton gyrofrequency, $\Omega_{p}=2\pi f_{cp}$–of different plasma modes as a function of wavevector, ($k_{\perp}$, $k_{\parallel}$) and equilibrium plasma parameters $\mathcal{P}$.
@@
Figure 4: Contour maps of PLUME damping (red) and growth (blue) rates as a function of heliocentric distance and $f_{obs}$, the frequencies where ion-scale waves that are observed by PSP during Encounters 1 and 2 for the positive and negative Doppler-shifted ICW cases.
For ICWs (panels (a,b)), $|\omega_{\text{plasma}}|$ is usually less than $k_{\parallel}V_{\text{SW}\parallel}$ at all $k_{\parallel}$ (see orange curve), and they have LH (RH) polarization upon positive (negative) Doppler shift which are plotted as red-solid (blue-dashed) curves. However, for FMWs (panels (c,d)), $|\omega_{\text{plasma}}|$ become strongly dispersive at small parallel scales gaining larger phase velocities, allowing them to exceed $k_{\parallel}V_{\text{SW}\parallel}$ in magnitude (see orange curve). While a positive Doppler-shifted FMW has RH polarization (blue-solid curve), a negative Doppler-shifted FMW can switch polarizations from LH (red-dashed curve) to RH (blue-dashed) moving from smaller to larger $k_{\parallel}$. Consequently, for these waves, the magnitude of $\omega_{sc}$ increases with $k_{\parallel}$, reaches a maximum, and then decreases before reversing sign (see the red-dashed curve). The Doppler shift routine employed in this work and the behaviors of ICWs and FMWs upon Doppler shifting are consistent with Bowen et al. (2020), Verniero et al. (2020), and McManus et al. (2024).
@@ 3.1 Observed ion-scale waves are consistent with transient ICWs
Figure 7: The ratio of the decay time, $\tau_{\text{decay}}$ to the nonlinear turbulent timescale at proton scale, $\tau_{nl}$ corresponding to the positive and negative Doppler-shifted ICW cases.
@@
In all intervals where ion-scale waves are observed, we evaluated $\gamma^{\text{max}\_\text{wave}}$, the damping rate corresponding to the frequency in $f_{obs}$ at which the fraction of wave energy is maximum. In all intervals where $\gamma^{\text{max}\_\text{wave}}<0$, that is the observed waves damp at the frequency where their energy is maximum, we estimated the timescales over which the waves decay by evaluating $\tau_{\text{decay}}=\frac{1}{\gamma^{\text{max}\_\text{wave}}}$. We then evaluated the turbulent non-linear timescale $\tau_{nl}=\frac{k_{\perp}^{-2/3}k_{0}^{-1/3}}{v_{A}}$ at proton scales ($k_{\perp}\rho_{p}\sim 1$) by assuming critical balance (Goldreich & Sridhar (1995)). Here $k_{0}=\frac{2\pi f_{\text{break}}}{V_{\text{SW}}}$ is the outer scale evaluated by subdividing Encounter 1 and 2 observations into 6-hour intervals and then calculating $f_{\text{break}}$, the spectral break frequency at which the transition from injection to inertial scales occurs. Figure 7 shows the ratio $\frac{\tau_{\text{decay}}}{\tau_{nl}}$ for all the positive and negative Doppler shifted ICW cases. For ICWs $\tau_{\text{decay}}$ and $\tau_{nl}$ are comparable and both have median values of 2-3 seconds. We further estimated the distance, $l_{prop}=\tau_{\text{decay}}v_{\text{phase}}^{\text{max}\_\text{wave}}$ that the observed waves propagate during the time their amplitudes damp by a factor of e = 2.718. Here $v_{\text{phase}}^{\text{max}\_\text{wave}}$ is the phase velocity at the frequency where the wave energy fraction is maximum. We find that the median value of $l_{prop}/\rho_{p}$ is $\sim 9$. Hence these ICWs are local and very short-lived.
@@
These waves are short-lived compared to both the cadence of the temperature anisotropy estimations used in this work (1 minute estimations from Huang et al., 2020) and the time length of the intervals considered in this analysis (15 minutes), and could be driven unstable by short intervals of temperature anisotropy beyond the appropriate threshold. The width of the observed wave frequencies in an interval could be a result of the aggregation of distinct waves from different regions with different plasma parameters and solar wind velocities–and hence Doppler shifts– compared to their interval-averaged values. The observed transient ICWs can be intermittently generated in unstable regions (Qudsi et al., 2020) and damp at nearby stable regions, mediating the local spatial transport of free energy associated with temperature anisotropy. To demonstrate this, we illustrate in Fig. 8 the ICW growth and damping rates calculated using both the 15 minute average plasma parameters compared against rates determined from the 1-minute values. For both intervals, the one-minute averages demonstrate some stable and unstable behavior, even though the 15 minute averages are either stable or unstable. Across all wavevectors, we see the damping rates for the stable waves determined by the 15-minute average are qualitatively similar to the 1-minute values, suggesting that the assumption that the waves damp at rates associated with the plasma parameters from the 15 minute averages is a reasonable statistical estimation. Therefore, we argue that the estimates of cyclotron damping rates in stable regions as well as the estimates of $f_{PLUME}^{wave\_growth}$, $\tau_{\text{decay}}$, and $l_{prop}$ that are evaluated using interval averaged temperature anisotropy values are reasonable.
@@ 4 Significant ICW dissipation onto young solar wind protons
Figure 9: Estimated dissipation rates of observed ion-scale waves via cyclotron damping (this work) and turbulence via Landau damping (SKM23) in all the analyzed intervals from PSP Encounters 1 and 2. The dissipation rates of positive (red) and negative (blue) Doppler shifted ICWs onto protons, the turbulent dissipation rates onto protons via Landau damping (green), and the inertial scale energy cascade rates evaluated using PSP observations by Bandyopadhyay et al. (2020) (blue star) are shown. Panel (a) shows moving means of all $Q_{p}$ values and Panel (b) shows moving medians of net damped intervals $Q_{p}<0$. Here dots (circles) represent magnitudes of energy damping (growing) rates i.e $-Q_{p}$ ($Q_{p}$).
@@
Upon evaluating dissipation rates of ion-scale waves by employing the routine described in section 2 for all intervals from PSP Encounters 1 and 2 between 0.166 and 0.25 au that were analyzed in SKM23, we find that ICWs can dissipate onto protons at significant rates that could explain the anisotropic heating of protons which enhances at closer distances to the Sun (Hellinger et al., 2011, 2013). Figure 9 shows the dissipation rates onto protons, $Q_{p}$ by cyclotron damping of ion-scale wave estimated in this work and Landau damping ($Q_{p,Landau}$, green) of turbulence estimated in SKM23. Here dots represent net damping rates and open circles represent net growth rates. Panels (a,b) show $Q_{p}$ via cyclotron damping of positive (red) and negative (blue) Doppler-shifted ICWs and the corresponding moving averages/medians (solid lines). The mean $Q_{p}$ values via cyclotron damping are usually an order of magnitude higher than $Q_{p}$ values via Landau damping and are in good agreement with the inertial scale energy cascade rates estimated at 36 and 54 solar radii by Bandyopadhyay et al. (2020) using PSP observations (blue stars). That the cyclotron damping rate is similar to the turbulent cascade rate is consistent with Bowen et al. (2024), who show that ion-scale waves are correlated with signatures of sub-ion-scale turbulence and suggest that the waves play an active role in dissipating turbulence. Note that for $Q_{p}$ via cyclotron resonance the mean is an order of magnitude higher than the median while for $Q_{p}$ via Landau resonance the mean and median are similar. We infer that the proton cyclotron damping is intermittent and bursty, while the proton Landau damping is constant and continuous. We find that in 11.86% of the intervals where waves are observed, there is a net ICW energy growth rate (open circles in panel (a)). This fraction is lower than the fraction of intervals (38%) where the interval averaged values of $\frac{T_{\perp,p}}{T_{\parallel,p}}$ determine a growing wave at any wavevector. This discrepancy arises because ion-scale waves are observed over a wide effective $k_{\parallel}$ range, while the instability due to temperature anisotropy arise over a narrower $k_{\parallel}$ band (as seen in Figure 4). Integrating the growth and damping rates over the observed scales often results in a net damping rate.
@@
We define a total possible heating rate due to cyclotron damping onto protons, $Q_{p,cyc}$, as the sum of the red and blue dots in panel (a). In intervals where both LH and RH polarized waves are observed, two distinct oppositely propagating ICWs can be positive and negatively Doppler shifted to independently produce both the observed polarizations, and there is no double counting in the evaluation of $Q_{p,cyc}$ by adding heating rates of both polarizations.
@@
Figure 10: Dominant heating of protons: (a) $Q_{p}/(Q_{p}+Q_{e})$ evaluated by calculating the moving averages of $Q_{p}$ and $Q_{e}$ by considering proton heating via Landau resonance (light green dashed), cyclotron resonance (medium green dash-dotted) and both the resonances (dark green solid) in intervals where waves are observed are shown (green shaded in Table 2). Further $Q_{p}/(Q_{p}+Q_{e})$ evaluated due to Landau resonance alone (red) in intervals where waves are not observed (red shaded in Table 2) as functions of heliocentric distance are shown. The fit of $Q_{p}/(Q_{p}+Q_{e})$ along with error bars estimated by Bandyopadhyay et al. (2023) is shown (cyan) for comparison. Additionally, the moving medians of (b) wave energy and (c) $\beta_{p}$ as functions of heliocentric distance are shown.

### v3

@@ 2.1 Plasma and Magnetic Field Data
We use Level 2 magnetic field data (in spacecraft frame RTN coordinates) from the Flux Gate Magnetometer (MAG) component of the PSP/FIELDS instrument suite, which makes in situ observations of electromagnetic fields (Bale et al., 2016). While the sampling rate of the MAG varies between $\sim$73.24 Sa $s^{-1}$, 146.48 Sa $s^{-1}$, and 292.96 Sa $s^{-1}$ during Encounter 1, the magnetic field is sampled at a constant rate of 146.48 Sa $s^{-1}$ during Encounter 2. Measurements of the thermal plasma come from the PSP/SWEAP instrument suite (Kasper et al., 2016). Huang et al. (2020) and Halekas et al. (2020) have estimated parallel and perpendicular temperatures (with respect to the direction of magnetic field) of protons ($T_{\parallel,p}$ and $T_{\perp,p}$) and electrons ($T_{\parallel,e}$ and $T_{\perp,e}$), respectively by applying bi-Maxwellian fits to velocity distribution functions (VDFs) of protons evaluated from SWEAP/SPC (Case et al., 2020) and electrons evaluated from SWEAP/SPANe (Whittlesey et al., 2020) measurements. We use plasma parameters evaluated by averaging over SPC and SPANe observations as inputs into the PLUME solutions.
@@
Figure 2: Overview of methodology to estimate ion-scale wave dissipation in the PSP interval (considered in Figure 1) on 10th November 2018 17:53:58 - 18:08:53 where both RH and LH polarized waves are observed: (a) The spacecraft-frame polarization spectrum, $\sigma(\omega,t)$, (b) the frequency ranges where LH and RH polarized waves are observed ($\omega_{\text{obs}}^{LH/RH}$, red and blue dashed-dotted, respectively), and (c) their corresponding energy spectra (red and blue) along with the total energy spectrum (black). (d, e) The damping rates of ICWs and FMWs onto protons and electrons, respectively as functions of spacecraft-frame frequencies that are evaluated using PLUME for all possible Doppler shifts.
@@ 2.3 Removing non-wave Signatures and Identifying Ion-scale Wave Frequencies in the Spacecraft Frame
In SKM23, we observed significant energy in ion-scale waves in the first two PSP encounters, as has been previously reported, e.g., Bowen et al. (2020b), Verniero et al. (2020), Zhao et al. (2021), and Liu et al. (2023). Panel (a) of Figure 2 shows $\sigma(s,t)$ for the same example PSP 15-minute interval on November 10, 2018, where ion-scale waves are observed. Persistent high values of $|\sigma|$ associated with these waves are observed over frequency bands near the proton cyclotron frequency, $f_{cp}=\frac{q_{p}B}{2\pi m_{p}}$, and last for times varying from several seconds to a few hours. Here $q_{p}$ is the charge of the proton. In the absence of ion-scale waves, non-persistent random $\sigma$ values corresponding to the background turbulent fluctuations are observed over timescales of $\sim$ 1/frequency, i.e. the time resolution of the wavelet transform. By employing a time-smoothing routine on $\sigma(s,t)$, we identify the frequency band, $f^{(RH/LH)}_{\text{obs}}$ where the ion-scale waves are observed (shown in panel (b) of Figure 2) and evaluate their energy spectra, $|\tilde{B}(f)|^{2^{LH/RH}}_{wave\_psp}$ (shown in panel (c) of Figure 2) as described in Appendix A. Here $\omega^{(RH/LH)}_{\text{obs}}=2\pi f^{(RH/LH)}_{\text{obs}}$ is the angular frequency band where the waves are observed.
@@ 2.4 Calculating Damping Rates at $\omega_{\text{obs}}$
The four MHD modes - the Alfvén, fast and slow magnetosonic waves, and the entropy mode have kinetic counterparts in a collisionless plasma that can be relatively lightly damped (see Klein (2013) for an overview). The assumption of symmetric VDFs produces degenerate solutions for the three propagating modes, and therefore the damping rates of the forward and backward solutions are identical. The impact of the mode propagation direction is considered in the calculation of the spacecraft to plasma frame conversion, see Eqn. 2. At small parallel scales, Alfvén waves extend to ICWs and FMWs extend to whistler waves. The PLUME dispersion solver (Klein & Howes, 2015) calculates the hot linear plasma dispersion relation $|D(\omega,\gamma;\mathbf{k},\mathcal{P})|=0$ for bi-Maxwellian VDFs of ions and electrons in a uniform magnetized environment. The solutions of the dispersion relation yield dimensionless angular frequencies and damping rates $(\omega/\Omega_{p},\gamma/\Omega_{p})_{\texttt{PLUME}}$ in the plasma frame–which are normalized to physical units via the observed proton gyrofrequency, $\Omega_{p}=2\pi f_{cp}$–of different plasma modes as a function of wavevector, ($k_{\perp}$, $k_{\parallel}$) and equilibrium plasma parameters $\mathcal{P}$.
@@
Figure 4: Contour maps of PLUME damping (red) and growth (blue) rates as a function of heliocentric distance and $f_{obs}$, the frequencies where ion-scale waves are observed by PSP during Encounters 1 and 2 for the positive and negative Doppler-shifted ICW cases.
For ICWs (panels (a,b)), $|\omega_{\text{plasma}}|$ is usually less than $k_{\parallel}V_{\text{SW}\parallel}$ at all $k_{\parallel}$ (see orange curve), and they have LH (RH) polarization upon positive (negative) Doppler shift which is plotted as red-solid (blue-dashed) curve. However, for FMWs (panels (c,d)), $|\omega_{\text{plasma}}|$ become strongly dispersive at small parallel scales gaining larger phase velocities, allowing them to exceed $k_{\parallel}V_{\text{SW}\parallel}$ in magnitude (see orange curve). While a positive Doppler-shifted FMW has RH polarization (blue-solid curve), a negative Doppler-shifted FMW can switch polarizations from LH (red-dashed curve) to RH (blue-dashed) moving from smaller to larger $k_{\parallel}$. Consequently, for these waves, the magnitude of $\omega_{sc}$ increases with $k_{\parallel}$, reaches a maximum, and then decreases before reversing sign (see the red-dashed curve). The Doppler shift routine employed in this work and the behaviors of ICWs and FMWs upon Doppler shifting are consistent with Bowen et al. (2020), Verniero et al. (2020), and McManus et al. (2024).
@@ 3.1 Observed ion-scale waves are consistent with transient ICWs
Figure 7: The ratio of the decay time, $\tau_{\text{decay}}$ to the nonlinear turbulent timescale at proton scale, $\tau_{nl}$ corresponding to the positively and negatively Doppler-shifted ICW cases.
@@
In all intervals where ion-scale waves are observed, we evaluated $\gamma^{\text{max}\_\text{wave}}$, the damping rate corresponding to the frequency in $f_{obs}$ at which the fraction of wave energy is maximum. In all intervals where $\gamma^{\text{max}\_\text{wave}}<0$, i.e., the observed waves damp at the frequency where their energy is maximum, we estimated the timescales over which the waves decay by evaluating $\tau_{\text{decay}}=\frac{1}{-\gamma^{\text{max}\_\text{wave}}}$. We then evaluated the turbulent non-linear timescale $\tau_{nl}=\frac{k_{\perp}^{-2/3}k_{0}^{-1/3}}{v_{A}}$ at proton scales ($k_{\perp}\rho_{p}\sim 1$) by assuming critical balance (Goldreich & Sridhar (1995)). Here $k_{0}=\frac{2\pi f_{\text{break}}}{V_{\text{SW}}}$ is the outer scale evaluated by subdividing Encounter 1 and 2 observations into 6-hour intervals and then calculating $f_{\text{break}}$, the spectral break frequency at which the transition from injection to inertial scales occurs. Figure 7 shows the ratio $\frac{\tau_{\text{decay}}}{\tau_{nl}}$ for all the positively and negatively Doppler shifted ICW cases. For ICWs $\tau_{\text{decay}}$ has a median value of 7.6 seconds and is comparable to $\tau_{nl}$. We further estimated the distance, $l_{prop}=\tau_{\text{decay}}v_{\text{phase}}^{\text{max}\_\text{wave}}$ that the observed waves propagate during the time their amplitudes damp by a factor of e = 2.718. Here $v_{\text{phase}}^{\text{max}\_\text{wave}}$ is the phase velocity at the frequency where the wave energy fraction is maximum. We find that the median value of $l_{prop}/\rho_{p}$ is $\sim$60. Hence these ICWs are local and very short-lived.
@@
These waves are short-lived compared to both the cadence of the temperature anisotropy estimations used in this work (1 minute estimations from Huang et al., 2020) and the time length of the intervals considered in this analysis (15 minutes), and could be driven unstable by short intervals with temperature anisotropy beyond the appropriate threshold. The width of the observed wave frequencies in an interval could be a result of the aggregation of distinct waves from different regions with different plasma parameters and solar wind velocities–and hence Doppler shifts– compared to their interval-averaged values. The observed transient ICWs can be intermittently generated in unstable regions (Qudsi et al., 2020) and damp at nearby stable regions, mediating the local spatial transport of free energy associated with temperature anisotropy. To demonstrate this, we illustrate in Fig. 8 the ICW growth and damping rates calculated using both the 15 minute average plasma parameters compared against rates determined from the 1-minute values. For both intervals, the one-minute averages demonstrate some stable and unstable behavior, even though the 15 minute averages are either stable or unstable. Across all wavevectors, we see the damping rates for the stable waves determined by the 15-minute average are qualitatively similar to the 1-minute values, suggesting that the assumption that the waves damp at rates associated with the plasma parameters from the 15 minute averages is a reasonable statistical estimation. Therefore, we argue that the estimates of cyclotron damping rates in stable regions as well as the estimates of $f_{PLUME}^{wave\_growth}$, $\tau_{\text{decay}}$, and $l_{prop}$ that are evaluated using interval averaged temperature anisotropy values are reasonable.
@@ 4 Significant ICW dissipation onto young solar wind protons
Figure 9: Estimated dissipation rates of observed ion-scale waves via cyclotron damping (this work) and turbulence via Landau damping (SKM23) in all the analyzed intervals from PSP Encounters 1 and 2. The dissipation rates of positively (red) and negatively (blue) Doppler shifted ICWs onto protons, the turbulent dissipation rates onto protons via Landau damping (green), and the inertial scale energy cascade rates evaluated using PSP observations by Bandyopadhyay et al. (2020) (blue star) are shown. Panel (a) shows moving means of all $Q_{p}$ values and Panel (b) shows moving medians of net damped intervals $Q_{p}<0$. Here dots (circles) represent magnitudes of energy damping (growing) rates i.e $-Q_{p}$ ($Q_{p}$).
@@
Upon evaluating dissipation rates of ion-scale waves by employing the routine described in Section 2 for all intervals from PSP Encounters 1 and 2 between 0.166 and 0.25 au that were analyzed in SKM23, we find that ICWs can dissipate onto protons at significant rates that could explain the anisotropic heating of protons which enhances at closer distances to the Sun (Hellinger et al., 2011, 2013). Figure 9 shows the dissipation rates onto protons, $Q_{p}$ by cyclotron damping of ion-scale wave estimated in this work and Landau damping ($Q_{p,Landau}$, green) of turbulence estimated in SKM23. Here dots represent net damping rates and open circles represent net growth rates. Panels (a,b) show $Q_{p}$ via cyclotron damping of positively (red) and negatively (blue) Doppler-shifted ICWs and the corresponding moving averages/medians (solid lines). The mean $Q_{p}$ values via cyclotron damping are usually an order of magnitude higher than $Q_{p}$ values via Landau damping and are in good agreement with the inertial scale energy cascade rates estimated at 36 and 54 solar radii by Bandyopadhyay et al. (2020) using PSP observations (blue stars). That the cyclotron damping rate is similar to the turbulent cascade rate is consistent with Bowen et al. (2024), who show that ion-scale waves are correlated with signatures of sub-ion-scale turbulence and suggest that the waves play an active role in dissipating turbulence. Note that for $Q_{p}$ via cyclotron resonance the mean is an order of magnitude higher than the median while for $Q_{p}$ via Landau resonance the mean and median are similar. We infer that the proton cyclotron damping is intermittent and bursty, while the proton Landau damping is constant and continuous. We find that in 12.74% of the intervals where waves are observed, there is a net ICW energy growth rate (open circles in panel (a)). This fraction is lower than the fraction of intervals (38%) where the interval averaged values of $\frac{T_{\perp,p}}{T_{\parallel,p}}$ determine a growing wave at any wavevector. This discrepancy arises because ion-scale waves are observed over a wide effective $k_{\parallel}$ range, while the instabilities due to temperature anisotropy arise over a narrower $k_{\parallel}$ band (as seen in Figure 4). Integrating the growth and damping rates over the observed scales often results in a net damping rate.
@@
We define a total possible heating rate due to cyclotron damping onto protons, $Q_{p,cyc}$, as the sum of the red and blue dots in panel (a). In intervals where both LH and RH polarized waves are observed, two distinct oppositely propagating ICWs can be positively and negatively Doppler shifted to independently produce both the observed polarizations, and there is no double counting in the evaluation of $Q_{p,cyc}$ by adding heating rates of both polarizations.
@@
Figure 10: Predicted proton-to-total heating rates $Q_{p}/(Q_{p}+Q_{e})$ as a function of heliospheric distance evaluated by calculating the moving averages of $Q_{p}$ and $Q_{e}$. We separately consider proton heating via the Landau resonance (light green dashed), the cyclotron resonance (medium green dashed-dotted) and both resonances (dark green solid) in intervals where waves are observed (green shaded in Table 2). We also show $Q_{p}/(Q_{p}+Q_{e})$ evaluated from Landau damping alone (red) in intervals where waves are not observed (red shaded in Table 2). The fit of $Q_{p}/(Q_{p}+Q_{e})$ along with error bars estimated by Bandyopadhyay et al. (2023) is shown (cyan) for comparison.